OpenStudy (anonymous):

x^2 +3x -7=0 Exact solutions x= _ Approximate solutions to 3 decimal places x=_

OpenStudy (anonymous):

Since it doesnt look like it can be factored, you will have to resort to using the quadratic formula: \[x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\] \[ax^2+bx+c\]\[x^2+3x-7\] so it looks like a = 1, b = 3, and c = -7

OpenStudy (anonymous):

plugging in and solving we get: \[x = \frac{-(3)\pm \sqrt{(3)^2-4(1)(-7)}}{2(1)}\]\[x = \frac{-3\pm \sqrt{9+28}}{2} = \frac{-3\pm \sqrt{37}}{2}\] cant be simplified anymore than that, those are the exact answers (one with the plus, and one with the minus)

OpenStudy (anonymous):

for the approximate answers, you would need to plug it into a calculator, which unfortunately i dont have on me at the moment. sry =/

OpenStudy (anonymous):

Thanks so much :)